# A Generalized Nonlinear extension of Quantum Mechanics

### v1

@@ I Introduction of the model
In a previous paper cf23 , we proposed a non-linear extension of quantum mechanics and studied some of its consequences. Our motivations were various: partly, in the spirit of Weinberg Weinberg89 and Kaplan and Rajendran Kaplan , we simply wanted to see what a particular non-linear extension of quantum mechanics might look like, and to deduce some of its consequences.
But we also had in mind the question of whether, as others have speculated Penrose Carlip , the union of gravity with quantum mechanics is best achieved not by producing a fully quantum version of general relativity, as is contemplated for example in string theory or loop quantum gravity, but by some other means, such as the coupling of quantum mechanics to a classical version of gravity, as in the work of Oppenheim and collaborators Oppenheim , or indeed by the modification of quantum mechanics itself.
In the model we introduced in cf23 ; cf24 , a given system is described by 2 state vectors, $|\psi>$ and $|\phi>$, in Hilbert space, in contrast to the single state vector to which ordinary quantum mechanics limits itself. In cf23 an exploration of the extra information provided by the pair of state vectors was not attempted. The model had the advantage of simplicity and was completely integrable, in the sense that the equations of motion determined the time dependence of the inner products $\psi|\psi>$, $\phi|\psi>$, and $\phi|\phi>$ explicitly in terms of a small number of integration constants, and one could then go on to determine the modification of the time dependence of the state vectors themselves due to the non-linearities. The non-linearity in that model is proportional to the inner product $\gamma=<\phi|\psi>$. One finds that $\gamma\rightarrow 0$ for large times, so that the effect of the non-linearity vanishes in the far future and the far past. In this work we seek a generalization that will allow for a broader range of dynamical possibilities. If we define
@@
$\mathcal{M=}\begin{pmatrix}0&g<\phi|\psi>\\ g^{*}<\psi|\phi>&0\\ \end{pmatrix}$ (4)
@@
$\mathcal{M}_{0}=\mathcal{M}_{0}^{\dagger}=\begin{pmatrix}\alpha_{1}<\psi|\psi>+\alpha_{2}<\phi|\phi>&g<\phi|\psi>\\ g^{*}<\psi|\phi>&\beta_{1}<\psi|\psi>+\beta_{2}<\phi|\phi>\\ \end{pmatrix}$ (7)
@@ I.1 The $\tau$- $\delta$ problem
We then have that $\tau$ and $\delta$ obey the following differential equations:
@@ II.1 The case p=1p=-1
$s=-\frac{i~F\left(\frac{i\kappa}{2}|\frac{8\sqrt{c}}{4\sqrt{c}-N^{2}}\right)}{\sqrt{N^{2}-4.\sqrt{c}}}$ (30)
@@
$\kappa=-2i~\text{am}\left(i\sqrt{N^{2}-4\sqrt{c}}s|\frac{8\sqrt{c}}{4\sqrt{c}-N^{2}}\right)$ (31)
which is shown in Fig.( 1) for $c=4,N^{2}=25$
@@
When $N^{2}=25.c=4$ we get the behavior for $\tau(s)$ shown in Fig. (2) .
@@ III Properties of the density matrix and the trajectory function
In cf23 we defined the density matrix
@@ III.1 Trajectory function for the simple solution
where ${\overrightarrow{X}}_{0}$, $\overrightarrow{V}$ and $\overrightarrow{W}$ are constant vectors depending on the values we choose for $A\left|X_{i}\right|A>,\ <A\left|X_{i}\right|B>$ and $B\left|X_{i}\right|B>$ . Thus, the particle moves in the plane spanned by $\overrightarrow{V}$ and $\overrightarrow{W}$. Calling the coordinates in this plane $X_{1}^{{}^{\prime}}$ and $X_{2}^{{}^{\prime}}$, we can orient the axes so that the equation for the orbit can be written in the form
@@ IV Conclusions and outlook
In a previous paper cf23 , we began an exploration of a particular form of non-linear extension of quantum mechanics. The extension could be applied to any underlying dynamical system, and featured the introduction of two state vectors instead of the usual single state vector in ordinary quantum mechanics. The model was simple, and the effects of the non-linearity could be determined completely, although the meaning of the second state vector remained uncertain. An analogy was noted with the two-state formalism of Aharonov and Vaidman Aharonov-Vaidman which is a time-symmetric formulation of ordinary quantum mechanics, based on earlier work by Aharonov, Bergmann and Lebowitz Aharonov64 . Because of our model’s simplicity, the dynamics it incorporated was limited. In particular, the effects of the non-linearity persisted for only a finite duration, decaying exponentially for large times. In addition, the natural extension of the Born rule to the model failed to encompass any of the non-linear behavior. Therefore, in this paper, we have constructed an expanded version, finding the most general such model consistent with three defining principles that were characteristic of the earlier work. Although we are not yet able to solve this new model completely, we can already see, by studying some simple examples, that the time dependence of the state vectors is much richer than that allowed by the earlier model. In particular, closed orbits exist for a range of parameters, which was not possible when the non-linearity only persisted for a finite time.

### v2

@@ I Introduction of the model
In a previous paper cf23 , we proposed a non-linear extension of quantum mechanics and studied some of its consequences. Our motivations were various: partly, in the spirit of Birula , Weinberg89 and Kaplan , we simply wanted to see what a particular non-linear extension of quantum mechanics might look like, and to deduce some of its consequences.
But we also had in mind the question of whether, as others have speculated Penrose Carlip Pang , the union of gravity with quantum mechanics is best achieved not by producing a fully quantum version of general relativity, as is contemplated for example in string theory or loop quantum gravity, but by some other means, such as the coupling of quantum mechanics to a classical version of gravity, as in the work of Oppenheim and collaborators Oppenheim , or indeed by the modification of quantum mechanics itself.
In the model we introduced in cf23 ; cf24 , a given system is described by 2 state vectors, $|\psi>$ and $|\phi>$, in Hilbert space, in contrast to the single state vector to which ordinary quantum mechanics limits itself. Our procedure is heuristic, that is, we do not impose a prior rationale for introducing two state vectors but instead look ahead to what the consequences are. In that sense, we follow the grand tradition of quantum mechanics, in which everybody knows how to calculate, but nobody knows what it all really means. Our approach affords us the opportunity to generalize quantum mechanics in what we consider to be an interesting way. In our first paper, we presented the minimal realization of this idea, which, surprisingly, allowed us to completely determine the effects of the non-linearity independently of the dynamics of the underlying quantum mechanical system. This simplicity, however, overly restricted the evolution possibilities. The nonlinearity in that model is proportional to the inner product $\gamma=<\phi|\psi>$. One finds that $\gamma\rightarrow 0$ for large times, so that the effect of the nonlinearity vanishes in the far future and the far past.
In this paper, we retain the same principles but allow for a greater range of dynamical effects at the cost of no longer being able to solve the system of equations exactly. Rather than speculate about the meaning of the second state vector, we study the effects it produces, thereby hoping to gain information on what meaning it might have. One line of investigation, which we began to consider in cf24 , is that the additional dynamics have something to do with gravity. In other words, perhaps gravity is, at least in part, a manifestation of the non-linear aspects of quantum mechanics that we have introduced. The fact that our non-linear modification is universal, i.e. independent of the Hamiltonian of the underlying system, may be relevant in this regard. In section III we define a trajectory function, which describes how a particle would move under the influence of the non-linear terms, and compute it in a simple case.
If we define
@@
$\mathcal{M=}\begin{pmatrix}0&g<\phi|\psi>\\ g^{*}<\psi|\phi>&0\\ \end{pmatrix}$. (4)
@@
The reason for condition (I) is that it encapsulates the philosophy behind our non-linear extension of quantum mechanics. There are infinitely many ways to try to extend quantum mechanics; the procedure we have chosen is to introduce a second state vector, and then to have the relevant inner products govern the dynamics of the non-linear extension. That is the content of condition (I). Condition (II) is important for two reasons: it is the appropriate generalization of the usual rule that the norm of the state vector remains constant, and it ensures that the evolution of the state vectors will remain bounded despite the presence of complex coupling coefficients. Condition (III) is a consequence of Occam’s razor. It is a symmetry that significantly restricts the number of allowed terms in the equations of motion.
@@
$\mathcal{M}_{0}=\mathcal{M}_{0}^{\dagger}=\begin{pmatrix}\alpha_{1}<\psi|\psi>+\alpha_{2}<\phi|\phi>&g<\phi|\psi>\\ g^{*}<\psi|\phi>&\beta_{1}<\psi|\psi>+\beta_{2}<\phi|\phi>\\ \end{pmatrix}$. (7)
@@ I.1 The $\tau$- $\delta$ problem
The following differential equations for $\tau$ and $\delta$ are a direct consequence of the equations of motion, Eq. (1.3):
@@ II.1 The case p=1p=-1
$s=-\frac{i~F\left(\frac{i\eta}{2}|\frac{8\sqrt{c}}{4\sqrt{c}-N^{2}}\right)}{\sqrt{N^{2}-4.\sqrt{c}}}$ (30)
@@
$\eta=-2i~\text{am}\left(i\sqrt{N^{2}-4\sqrt{c}}s|\frac{8\sqrt{c}}{4\sqrt{c}-N^{2}}\right)$ (31)
When $c=4,\eta_{0}=0$. so that $\kappa=\eta$. We plot $\kappa(s)$ in Fig.( 1) for $c=4,N^{2}=25$
@@
When $N^{2}=25$ and $c=4$ we get the behavior for $\tau(s)$ shown in Fig. (2) .
@@ III Properties of the density matrix and the trajectory function
In cf23 we defined the density matrix , with its trace properly normalized to 1,
@@
In conventional quantum mechanics, the deviation of the trace of $\rho^{2}$ from 1 is a measure of a lack of purity of the state. Here we see that this quantity is both generally less than one, and dependent on time. It is only equal to 1 if $c=0$, a limit in which $|\psi>$ and $|\phi>$ are proportional (because $\mathcal{S}=0$).
@@ III.1 Trajectory function for the simple solution
where ${\overrightarrow{X}}_{0}$, $\overrightarrow{V}$ and $\overrightarrow{W}$ are real constant vectors depending on the values we choose for $A\left|X_{i}\right|A>,\ <A\left|X_{i}\right|B>$ and $B\left|X_{i}\right|B>$ . Thus, the particle moves in the plane spanned by $\overrightarrow{V}$ and $\overrightarrow{W}$. Calling the coordinates in this plane $X_{1}^{{}^{\prime}}$ and $X_{2}^{{}^{\prime}}$, we can orient the axes so that the equation for the orbit can be written in the form
@@ IV Conclusions and outlook
In a previous paper cf23 , we began an exploration of a particular form of non-linear extension of quantum mechanics. The extension could be applied to any underlying dynamical system, and featured the introduction of two state vectors instead of the usual single state vector in ordinary quantum mechanics. The model was simple, and the effects of the non-linearity could be determined completely, although the meaning of the second state vector remained uncertain. An analogy was noted with the two-state formalism of Aharonov and Vaidman Aharonov-Vaidman which is a time-symmetric formulation of ordinary quantum mechanics, based on earlier work by Aharonov, Bergmann and Lebowitz Aharonov64 . We had found that our original model possessed a time-reversal symmetry, involving the interchange of $|\psi>$ and $|\phi>$. In the present extended model, this symmetry is apparently broken by the parameter $\lambda$, as we explain in the appendix.
Because of our previous model’s simplicity, the dynamics it incorporated was limited. In particular, the effects of the non-linearity persisted for only a finite duration, decaying exponentially for large times. In addition, the natural extension of the Born rule to the model failed to encompass any of the non-linear behavior. Therefore, in this paper, we have constructed an expanded version, finding the most general such model consistent with three defining principles that were characteristic of the earlier work. Although we are not yet able to solve this new model completely, we can already see, by studying some simple examples, that the time dependence of the state vectors is much richer than that allowed by the earlier model. In particular, closed orbits exist for a range of parameters, which was not possible when the non-linearity only persisted for a finite time. Previous attempts to make quantum mechanics non-linear, such as that of Weinberg, were shown to permit superluminal communication Gisin Polchinski . That is, if two widely separated observers, Alice and Bob, were each in possession of a piece of an entangled state, Alice’s measurement would instantaneously collapse the wave function describing Bob’s piece. In ordinary quantum mechanics, however, Bob could not deduce from the collapse what measurement Alice made. In many non-linear extensions, though, this is no longer true, so Bob can obtain information from Alice instantaneously, violating the so-called no-signaling condition. Traditionally, this has been taken to be a fatal disease. The rationale given in one paper Helou is: “The no-signaling condition states that one cannot send information faster than the speed of light, and is a cornerstone of the special theory of relativity. The community regards this condition as being inviolable.”
We do not yet know whether our model allows this kind of superluminal communication, but if so, we do not regard it as fatal, or even necessarily detrimental. The effect will undoubtedly be hard to observe, and its discovery would reveal an interesting property of nature. We stress that physics is not determined by a vote of “the community”, but rather by the results of reliable experiments.
@@
Appendix A Note on Time-reversal
Let $K$ be an anti-unitary operator that commutes with $H$ and define $K|A>=|\tilde{A}$> for any state $|A>$. Because $K$ is anti-unitary, we must keep track of the direction in which $K$ acts. We denote its direction of action with an arrow. We have $\vec{K}\vec{K}=1$, so
$B|A>=<B|\vec{K}\vec{K}|A>=<B|\overleftarrow{K}~\overrightarrow{K}|A>^{*}=<\tilde{A}|\tilde{B}$>. (104)
Now we define $|\psi^{\prime}(t)>=K|\phi(-t)>=|\tilde{\phi}(-t)>$ and $|\phi^{\prime}(t)>=K|\psi(-t)>=|\tilde{\psi}(-t)>$.
Hence we have
$<\psi^{\prime}(t)|\psi^{\prime}(t)>$ $=<\tilde{\phi(-t)}|\tilde{\phi}(-t)>=<\phi(-t)|\phi(-t)>$
$<\phi^{\prime}(t)|\phi^{\prime}(t)>$ $=<\tilde{\psi}(-t)|\tilde{\psi}(-t)>=<\psi(-t)|\psi(-t)>,$ (105)
and
$\phi^{\prime}(t)|\psi^{\prime}(t)>=<\tilde{\psi}(-t)|\tilde{\phi}(-t)>=<\phi(-t)|\psi(-t)>$ (106)
i.e. $\gamma^{\prime}(t)=\gamma(-t).$
We evaluate the equations of motion at $t\rightarrow-t$ setting the $\alpha_{i}$ and $\beta_{i}$ to zero for simplicity and obtain:
$i\frac{d|\psi(-t)>}{dt}=\left[H+i\mu<\phi(-t)|\phi(-t)>\right]~|\psi(-t)>+(g+\lambda)\gamma(-t)|\phi(-t)>$ (107)
and
$i\frac{d|\phi(-t)>}{dt}=\left[H-i\mu<\psi(-t)|\psi(-t)>\right]~|\phi(-t)>+(g^{*}-\lambda)\gamma^{*}(-t)|\psi(-t)>.$ (108)
Now we apply $K$, assuming $[K,H]=0$, and noting that $K$ will take the complex conjugate of all coefficients we then obtain”
$i\frac{d|\phi^{\prime}(t)>}{dt}=\left[H-i\mu<\psi^{\prime}(t)|\psi^{\prime}(t)>\right]~|\phi^{\prime}(t)>+(g^{*}+\lambda)\gamma^{\prime*}(t)|\psi^{\prime}(t)>$ (109)
and
$i\frac{d|\psi^{\prime}(t)>}{dt}=\left[H+i\mu<\phi^{\prime}(t)|\phi^{\prime}(t)>\right]~|\psi^{\prime}(t)>+(g-\lambda)\gamma^{\prime}(t)|\phi^{\prime}(t)>.$ (110)
We recognize these as the equations satisfied by $|\phi(t)>$ and $|\psi(t)>$, respectively, except for one thing: the sign of $\lambda$ . So, modulo the problem with $\lambda$ the system possesses a time-reversal symmetry. It is not clear at this point whether there is a more general definition of the time reversal transformation that will remove the problem with $\lambda$ , or whether this is a fundamental source of time-reversal symmetry violation in our model.